OpenStudy (karatechopper):

Domain of sqrt(x^2+4)

OpenStudy (solomonzelman):

\(\large\color{black}{ x^2+4 }\) is always a positive number.

OpenStudy (danjs):

The domain are the possible values for x that will produce a output y

OpenStudy (karatechopper):

Are you saying all real numbers?

OpenStudy (solomonzelman):

Yes, good job!

OpenStudy (solomonzelman):

\(\large\color{black}{ x^2\ge0 }\) ( for all real values of x.) that means that you will never get any imaginary values in this case, (no matter what x is).

OpenStudy (karatechopper):

Could you explain why?

OpenStudy (karatechopper):

Hm..its not quite clicking

OpenStudy (freckles):

0^2 5^2 (-5)^2 (6)^2 (-2523523)^2 are all either positive or 0 right?

OpenStudy (danjs):

If a square root function has a negative number underneath the root, the result is an imaginary number, you do not want that

OpenStudy (karatechopper):

Yes freckles.

OpenStudy (freckles):

so adding a positive number to any of those is still going to be positive

OpenStudy (karatechopper):

Correct Dan, elaborate more?

OpenStudy (karatechopper):

Well ok yes freckles, elaborate more please.

OpenStudy (karatechopper):

Are you all saying that when there is a positive number underneath a sq rt its always going to be all real numbers?

OpenStudy (freckles):

the domain is therefore all real numbers since x^2+4 is always positive

OpenStudy (danjs):

so you want x^2 + 4 >= 0 any x value squared , even if it is a negative number, will result in a positive value For example, (-2)^2 = +4

OpenStudy (solomonzelman):

yes, sure; For example if you had: \(\large\color{black}{ f(x)=\sqrt{x-4} }\) then you would know that when \(\large\color{black}{ x<4 }\) then you are taking the square root of a negative number. This is the only case where a square root can be undefined at x=c i.e. when the value of c (that you plug in for x, makes the inside of the square root a negative. In your case, you have: \(\large\color{black}{ f(x)=\sqrt{x^2+4} }\) and there you will never encounter a negative inside the root.

OpenStudy (freckles):

\[\sqrt{x-4} \text{ we have \to worry \because x-4 can be negative }\]

OpenStudy (freckles):

lol @SolomonZelman we thought of the same function

OpenStudy (solomonzelman):

yeah, that is weird and awesome!

OpenStudy (karatechopper):

Is it a rule that when you have imaginary in the sq rt its always going to be all real? If you can't fully solve it out, it's all real?

OpenStudy (danjs):



OpenStudy (solomonzelman):

the rule is that when you have a negative inside the square root (or any EVENth root) that it is then undefined.

OpenStudy (solomonzelman):

Like if you had: (again using,) \(\large\color{black}{ f(x)=\sqrt{x-4} }\) and plugged in 3 for x, you would be getting \(\large\color{black}{ f(x)=\sqrt{-1} }\) and that is undefined for any real value.

OpenStudy (karatechopper):

ok ok.

OpenStudy (karatechopper):

one more question

OpenStudy (solomonzelman):

I mean give me a real number a, so that: \(\large\color{black}{a \times a=-1 }\)

OpenStudy (solomonzelman):

not a possibility, right?

OpenStudy (solomonzelman):

sure, go ahead

OpenStudy (karatechopper):

I am doing compositions. g(f(x)) = sqrt(x+4)^2 =x+4 correct? Then I am doing the domain of it. I have to take the restrictions from both don't I? There was no restriction for g(x) it was all real and for f(x) it was x is greater than equal to 4. g(f(x)) is sqrt(x+4) so I got a domain of x is greater than equal to 4.

OpenStudy (solomonzelman):

you \(\large\color{black}{ f(x)=\sqrt{(x+4)^2} }\) is same as \(\large\color{black}{ f(x)=x+4 }\)

OpenStudy (solomonzelman):

and a polynomial is by definition continuous over \(\large\color{black}{ (- \infty,+\infty) }\).

OpenStudy (karatechopper):

yeah

OpenStudy (karatechopper):

wait what?!

OpenStudy (solomonzelman):

you are correct, and are there any restrictions, or not>?

OpenStudy (karatechopper):

crap.

OpenStudy (solomonzelman):

what?

OpenStudy (karatechopper):

lol

OpenStudy (solomonzelman):

you don't get something?

OpenStudy (karatechopper):

hold on. Lemme check this work real fast.